# Generalizable 3D Scene Reconstruction via Divide and Conquer from a Single View

### v1

@@ Abstract
Single-view 3D reconstruction is currently approached from two dominant perspectives: reconstruction of scenes with limited diversity using 3D data supervision or reconstruction of diverse singular objects using large image priors. However, real-world scenarios are far more complex and exceed the capabilities of these methods. We therefore propose a hybrid method following a divide-and-conquer strategy. We first process the scene holistically, extracting depth and semantic information, and then leverage a single-shot object-level method for the detailed reconstruction of individual components. By following a compositional processing approach, the overall framework achieves full reconstruction of complex 3D scenes from a single image. We purposely design our pipeline to be highly modular by carefully integrating specific procedures for each processing step, without requiring an end-to-end training of the whole system. This enables the pipeline to naturally improve as future methods can replace the individual modules. We demonstrate the reconstruction performance of our approach on both synthetic and real-world scenes, comparing favorable against prior works. Project page: https://andreeadogaru.github.io/Gen3DSR.
Keywords:
3D scene reconstruction single-view compositional
1 Introduction
Refer to caption Refer to caption Refer to caption Refer to caption
@@ 1 Introduction
Figure 1: Our method can reconstruct a scene in 3D from a single input view. The scene is semantically processed and partially occluded objects are completed. The individually reconstructed object instances are composed back into the scene using monocular depth guides. We represent each component with a triangle mesh, enabling downstream applications such as scene manipulation and editing.
Single-view 3D scene reconstruction refers to the problem of understanding and explaining all the visible components that assembled together create a 3D scene which closely reproduces the original 2D observation. The computer vision and graphics communities have long been interested in automating this task, yet its complexity still leaves room for many improvements [59, 30]. Successful single-view applications have been developed for specific purposes such as face reconstruction [14, 29], hair modeling [63], and many more. However, the 3D understanding from a single image task is far from solved in the case of larger scale problems such as indoor/outdoor scene reconstruction with multiple objects [54].
@@
Given the complexity of real-world scenes, reversing the process of image capturing in an end-to-end fashion would require a huge amount of data covering the variability of realistic environments. Therefore, many works solve a simplified version of the task by focusing on single objects or indoor rooms with a limited number of objects in which they predict the scene layouts containing 3D instance bounding boxes and optionally object meshes. Under these assumptions, most of the existent solutions rely on 3D scene geometry supervision from synthetic datasets. This class of methods usually struggles when applied to real-world images due to the domain gap. In contrast, we propose to tackle the single-view 3D scene reconstruction problem in a divide-and-conquer approach while building on the research advancement of related, simpler, tasks. In figure 1 we show that, following this approach, our pipeline is able to reconstruct multi-object scenes with unprecedented quality.
In the past few years, the field of computer vision has seen tremendous progress in solving particular tasks such as depth estimation and single-image 3D object reconstruction. It is the right time for these components to be assembled to solve the challenging task of full 3D scene reconstruction. We have identified the following sub-problems that together comprehensively explain a 3D scene and enable its reconstruction from a single input image: estimating the camera calibration, predicting the (metric) depth map, segmenting entities, detecting foreground instances, reconstructing the background, recovering the occluded parts of the individual objects (amodal completion), and reconstructing them. Designed as a disentangled framework, our method is open for incremental improvements and future enhanced components can be easily plugged in to boost the reconstruction performance of the entire system.
The main contributions of this paper are outlined as follows:
  • We build a compositional framework that comprehensively reconstructs a scene without requiring end-to-end training.
  • We develop a model for amodal completion and show how it can be used towards achieving full scene reconstruction.
  • We design the connecting links which enable integrating individually reconstructed 3D objects into the scene layout by exploiting single-view depth estimation.
  • We demonstrate through extensive quantitative and qualitative evaluations that the proposed modular approach exhibits generalizable full 3D scene reconstruction abilities.
@@ 2 Related Work
Starting with the pioneering work of Lawrence Roberts [51], the task of recovering 3D scene properties such as geometry, texture, and layout from a single 2D image has been extensively studied. Learning-based advancements have substantially progressed for this challenging task. However, existing methods still have limited understanding and struggle to faithfully reconstruct complex realistic scenes. In this section, we review the major lines of work and highlight the advantages of our approach that enable state-of-the-art results for realistic 3D scene reconstruction from a single view.
Feed-forward scene reconstruction.
@@
The direct regression conditioned on the input image used in many computer vision tasks can also be employed for 3D scene reconstruction. Provided the availability of large scene collections, such systems are trained end-to-end using 3D supervision. Most such works [10, 71, 9] rely on an encoder-decoder architecture that takes as input the image and predicts voxel grids containing scene properties such as geometry, semantics and instance labels. These methods have the advantage that by predicting the scene layout jointly with the containing objects, the object poses are intrinsically correct. However, these solutions suffer from resolution limitations, require large 3D data collections for training [15, 16] and usually do not generalize well to real-world scenes. In contrast, we use a modular framework that does not require end-to-end training or 3D supervision.
Factorized scene reconstruction.
@@
Following the formulation of Tulsiani et al. [60], the scene can be considered as composed of different factors including layout, object shape and pose that together make up a 3D scene representation. As this approach reconstructs the scene components individually, it is beneficial for downstream applications and the reconstruction system can be designed with specialized components for each type of factor. Methods following this paradigm initially reconstructed the objects in the scene as bounding boxes [20, 56, 13] or voxels [60, 36, 33], while later works use meshes [17, 45, 21, 34] or neural fields [37, 70, 69, 57]. Apart from the elementary factors of [60], more recent methods also include camera attributes [45, 21, 37, 20], textures [65, 8, 67], lightning [65, 67], or even material properties [65, 67]. Depending on the focus of the method, some works rely on a retrieval algorithm for identifying the object candidates [65, 21, 23, 2, 34] or regress their geometry [17, 37, 70, 57]. The initially predicted scene components are sometimes further fine-tuned using derivative-free optimization [23, 21, 7] or differentiable rendering [17, 67, 65, 69]. Our method also uses a compositional scene reconstruction framework. Particularly in our pipeline, each component is trained individually on extensive datasets. This strategy allows our proposed solution to transcend predefined (limited) classes of objects and reconstruct diverse scenes effectively.
Single-view scene understanding.
@@
The reconstruction of simple scenes composed of a single isolated object has seen numerous approaches along the years [18]. Nonetheless, enabled by the versatile diffusion-based models capable of generating realistic images [53, 47] and large collections of 3D objects [12, 11], there has recently been a surge of single-view 3D reconstruction methods [39, 38, 40, 58] capable of creating digital objects with unprecedented quality. Closely related to 3D scene reconstruction is monocular depth estimation [42], which predicts from a single image a 2.5D representation of the visible content. Recent methods focus on improving the generalizability of the estimator by following a training protocol employing large datasets [4, 66] or by fine-tuning a large image prior such as Stable Diffusion [53] for the depth estimation task [27]. A scene is also described by the camera used to capture the image [26, 35, 19] and by the separation between different instances [31, 49, 68]. Though the current state-of-the-art methods for each individual sub-task achieve robust results for real images, they only offer an incomplete explanation of the scene. Therefore, we compose them in an open framework capable of fully reconstructing complex real-world scenes from a single image.
@@ 3 Method
Figure 2: Method Overview: the input image is first analyzed collectively by an ensemble of state-of-the-art monocular models. Subsequently, the identified instances are individually processed, as elaborated in figure 3. The reconstructed objects, along with the modeled background, are composed into the final scene, rendered here and suitable to various applications.
@@ 3.1 Scene analysis
Our framework decouples the object-agnostic from the object-specific processing, pursuing a balance between the representational power and the generalization capabilities of the composing models. The image characteristics that can be best explained by perceiving the scene as a whole are the perspective properties and the depth. Additionally, the global context provides valuable guides for the recognition of the objects.
We rely on the Perspective Fields [26] formulation for estimating the camera calibration, $K_{img}$. Following the per-pixel parametrization using up-vectors and latitude, the representation can accommodate arbitrary camera models. In our pipeline we employ conventional camera parameters, i.e. field-of-view and principal point, which can be directly estimated from the Perspective Fields.
Monocular depth estimation is essential to our approach as it provides geometrical clues of scene layout, enabling the reconstruction of complex scenes that go beyond a single object. The performance of a depth estimator network is mostly determined by its capacity and training data. As we aim for a generalizable system, the single-image depth estimation method should perform well for a wide variety of natural images.
To fully explain the image, we employ entity segmentation [49, 50], which segments all semantically-meaningful entities without being constrained to a predefined set of classes. The unrestricted segmentation model yields robust dense predictions suitable for real-world scenarios. This step partitions each image $I$ into instances $\{M_{i}\}$ that can be individually reconstructed to compose the whole scene. Furthermore, we consider the natural separation of the instances in thing and stuff using a universal image segmentation model. This enables our method to tailor the reconstruction process for each group, leveraging their unique properties, as outlined in 3.2 and 3.3 respectively. To enhance downstream applications and text-conditioned reconstruction models, we harness the entire scene context to recognize and label the previously identified entities. Each segmented entity $M_{i}$ serves as a prompt for the OVSAM model [68], which, empowered by the recognition capabilities of the CLIP foundation model [25], outputs a corresponding instance label $L_{i}$.
@@ 3.2 Instance processing
Figure 3: Detailed Overview over the processing steps of each instance. An image is processed through the scene analysis part of our framework as described in figure 2. We add an object recognition information for diffusion guided completion to restore partially occluded objects. Then, we perform reconstruction and align the result back to the input view space.
@@
We consider an object $O_{i}$ to be a well-defined shape categorised as thing, identified by its entity mask $M_{i}$, with corresponding RGB-D region $I_{i}=I[M_{i}]$, $D_{i}=D[M_{i}]$, and a label $L_{i}$. The processing steps are illustrated in figure 3. We choose to reconstruct each instance individually in order to fully benefit from a view-conditioned 2D diffusion model $\mathcal{Z}$ [40, 11], trained using multi-view images of objects from large scale object collections [12, 11]. As the images used to train these models were rendered using a fixed predefined camera configuration $K_{crop}$, they do not generalize well to in-the-wild object crops. Therefore, we propose to reproject the object-associated pixels from the input image to the image space expected by the diffusion model.
To this end, we employ the previously estimated camera calibration $K_{img}$ and depth map $D_{i}$ to unproject the pixels into the view space as $P_{i}^{view}\in\mathbb{R}^{3}$. We then find the virtual camera that together with the desired intrinsics $K_{crop}$ closely match the observed image. Finally, we use the transformation to reproject the pixels to a crop $C_{i}$ that resembles the training domain of $\mathcal{Z}$.
@@
For simple scenes in which there is no occlusion between instances, the crop $C_{i}$ obtained using the equation above represents the full view of the object $O_{i}$. However, this is not the case for most real-world scenes, and directly feeding $C_{i}$ to the object reconstruction method $\mathcal{R}$ would result in an incomplete object. Therefore, we propose to recover the missing parts of $C_{i}$ by leveraging the image prior embodied by pre-trained large-scale diffusion models like Stable Diffusion. We approach the task named amodal completion which deals with recovering the shape and appearance of partially visible instances as an image-to-image translation problem. Specifically, we train a model to predict the view $\hat{C}_{i}$ of the full object $O_{i}$ conditioned on the object parts depicted in $C_{i}$ and the label $L_{i}$. Given the difficulty of collecting well-segmented training images guaranteed to contain complete objects, we generate a new synthetic dataset specifically for this task. We render synthetic objects from a large collection and then obtain the conditioning images by masking out parts of the object with a randomly overlaying silhouette of another object. We use this dataset to fine-tune Stable Diffusion, following the methodology of InstructPix2Pix [5], and concatenate the encoded conditioning image to the noisy latent. The weights added to the base network are initialized with zero, while the rest are taken from the pre-trained model to benefit from the learned image prior.
@@
The complete crop $\hat{C}_{i}$ can now be used as input for the single-image 3D object reconstruction method $\mathcal{R}$. As our pipeline is not restricted to a specific model, we describe their common structure. Using a view-conditioned diffusion prior such as Zero-1-to-3 [40, 11] enables these models to reconstruct a wide-range of objects from a single view without the need for 3D training supervision. The object is reconstructed using differentiable 3D representations, e.g. neural fields [43, 44], 3D Gaussians [28], that are either directly fitted to multi-view images generated by the diffusion prior, or by optimizing a Score Distillation Sampling-based loss [48] against the diffusion prior.
Then, a polygonal mesh $R_{i}^{obj}$ aligned to the input crop $\hat{C}_{i}$ is extracted from the neural field representation for use in downstream applications. The obtained mesh can optionally be further fine-tuned to refine the texture of the reconstructed object.
@@
The reconstructed instance in the object space $R_{i}^{obj}$ is transformed to the view space with respect to the inverse transformation determined by the virtual camera used for projection. The obtained mesh $R_{i}^{view}$ is aligned to the object points $P_{i}$ up to an unknown scale factor $s_{i}$; that is because $\mathcal{R}$ reconstructs objects at an arbitrary scale. We estimate $s_{i}$ as the scale factor that minimizes the distance between the visible points in $R_{i}^{view}$ and $P_{i}^{view}$.

### v2

# Gen3DSR: Generalizable 3D Scene Reconstruction via Divide and Conquer from a Single View

@@ Abstract
Single-view 3D reconstruction is currently approached from two dominant perspectives: reconstruction of scenes with limited diversity using 3D data supervision or reconstruction of diverse singular objects using large image priors. However, real-world scenarios are far more complex and exceed the capabilities of these methods. We therefore propose a hybrid method following a divide-and-conquer strategy. We first process the scene holistically, extracting depth and semantic information, and then leverage an object-level method for the detailed reconstruction of individual components. By splitting the problem into simpler tasks, our system is able to generalize to various types of scenes without retraining or fine-tuning. We purposely design our pipeline to be highly modular with independent, self-contained modules, to avoid the need for end-to-end training of the whole system. This enables the pipeline to naturally improve as future methods can replace the individual modules. We demonstrate the reconstruction performance of our approach on both synthetic and real-world scenes, comparing favorable against prior works. Project page: https://andreeadogaru.github.io/Gen3DSR
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
@@
Figure 1: Our method can reconstruct a 3D scene from a single view. We identify distinct objects, address their occlusions through amodal completion, and reconstruct them individually. The resulting 3D objects are composed into the scene using monocular depth guides. Each component is reconstructed as a triangle mesh, enabling downstream applications such as scene manipulation and editing.
1 Introduction
Single-view 3D scene reconstruction refers to the problem of understanding and explaining all the visible components that assembled together create a 3D scene which closely reproduces the original 2D observation. The computer vision and graphics communities have long been interested in automating this task, yet its complexity still leaves room for many improvements [71, 37]. Successful single-view applications have been developed for specific purposes such as face reconstruction [17, 36] and hair modeling [75]. However, 3D understanding from a single image is far from solved in the case of larger scale problems such as indoor/outdoor scene reconstruction with multiple objects [65].
@@ 1 Introduction
Given the complexity of real-world scenes, reversing the process of image capturing in an end-to-end fashion would require a huge amount of data covering the variability of realistic environments. Therefore, many works solve a simplified version of the task by focusing on single objects or indoor rooms with a limited number of object classes. Under these assumptions, most of the existent solutions rely on 3D scene geometry supervision from synthetic datasets. This class of methods usually struggles when applied to real-world images due to the domain gap and limited diversity in existing datasets. In contrast, we propose to tackle the single-view 3D scene reconstruction problem in a divide-and-conquer approach while building on the advances in related, simpler tasks. In Figure 1 we show that, following this approach, our pipeline is able to reconstruct multi-object scenes with unprecedented quality.
In the past few years, the field of computer vision has seen tremendous progress in solving particular tasks such as depth estimation and single-image 3D object reconstruction. It is the right time for these components to be assembled to solve the challenging task of full 3D scene reconstruction. We have identified the following sub-problems that together comprehensively explain a 3D scene and enable its reconstruction from a single input image: estimating the camera calibration, predicting the (metric) depth map, segmenting entities, detecting foreground instances, reconstructing the background, recovering the occluded parts of the individual objects (amodal completion), and reconstructing them. Our disentangled framework is open for incremental improvements and future enhanced modules can be easily plugged in to boost the reconstruction performance of the entire system. The main contributions of this paper are as follows:
  • We design a compositional framework and the corresponding abstractions, enabling scene-level 3D reconstruction without end-to-end training.
  • We build a model for amodal completion and show how it can be used towards achieving full scene reconstruction.
  • We develop the connecting links for integrating individually reconstructed 3D objects into the scene layout by exploiting single-view depth estimation.
  • We achieve an unmatched level of generalizability for real-world single-view 3D scene reconstruction, which we demonstrate through extensive evaluations.
@@ 2 Related Work
Starting with the pioneering work of Lawrence Roberts [62], the task of recovering 3D scene properties such as geometry, texture, and layout from a single 2D image has been extensively studied. Learning-based advancements led to substantial progress in this challenging task. However, existing methods still have limited understanding and struggle to faithfully reconstruct complex realistic scenes. In this section, we review the major lines of work and highlight the advantages of our approach that enable state-of-the-art results for 3D scene reconstruction from a single view.
@@
Feed-forward scene reconstruction. The direct regression conditioned on the input image used in many computer vision tasks can also be employed for 3D scene reconstruction. Provided the availability of large scene collections, such systems are trained end-to-end using 3D supervision. Most such works [12, 85, 11] rely on an encoder-decoder architecture that takes as input the image and predicts voxel grids containing scene properties such as geometry, semantics and instance labels. These methods have the advantage that by predicting the scene layout jointly with the containing objects, the object poses are intrinsically correct. However, these solutions suffer from resolution limitations, require large 3D data collections for training [19, 20] and usually do not generalize well to real-world scenes. In contrast, we use a modular framework that does not require end-to-end training or 3D supervision.
@@
Factorized scene reconstruction. Following the formulation of Tulsiani et al. [72], the scene can be considered as composed of different factors, including layout, object shape and pose, that together make up a 3D scene representation. As this approach reconstructs the scene components individually, it is beneficial for downstream applications and the reconstruction system can be designed with specialized components for each type of factor. Methods following this paradigm initially reconstructed the objects in the scene as bounding boxes [27, 67, 16] or voxels [72, 43, 40], while later works use meshes [22, 55, 28, 41] or neural fields [45, 84, 83, 68]. Apart from the elementary factors of [72], more recent methods also include camera attributes [55, 28, 45, 27], textures [79, 10, 81], lightning [79, 81], or even material properties [79, 81]. Depending on the focus of the method, some works rely on a retrieval algorithm for identifying the object candidates [79, 28, 30, 3, 41] or regress their geometry [22, 45, 84, 68]. The initially predicted scene components are sometimes further fine-tuned using derivative-free optimization [30, 28, 9] or differentiable rendering [22, 81, 79, 83]. Our method also uses a compositional scene reconstruction framework; in our pipeline, each module is trained individually on specific datasets. This strategy allows our proposed solution to transcend predefined (limited) classes of objects and reconstruct diverse scenes effectively.
Figure 2: Method Overview: the input image is first analyzed collectively by an ensemble of state-of-the-art monocular models. Subsequently, the identified instances are individually processed, as elaborated in Figure 3. The reconstructed objects, along with the modeled background, are composed into the final scene, which can then be used in various applications.
@@
Single-view scene understanding. The reconstruction of simple scenes composed of a single isolated object has seen numerous approaches along the years [24]. Nonetheless, enabled by the versatile diffusion-based models capable of generating realistic images [64, 57] and large collections of 3D objects [13, 14], there has recently been a surge of single-view 3D reconstruction methods [46, 47, 48, 69] capable of creating digital objects with unprecedented quality. Closely related to 3D scene reconstruction is monocular depth estimation [51], which predicts from a single image a 2.5D representation of the visible content. Recent methods focus on improving the generalizability of the estimator by following a training protocol employing large datasets [5, 80] or by fine-tuning a large image prior such as Stable Diffusion [64] for the depth estimation task [34]. A scene is also described by the camera used to capture the image [33, 42, 25] and by the separation between different instances [38, 60, 82]. Though the current state-of-the-art methods for each individual sub-task achieve robust results for real images, they only offer an incomplete explanation of the scene. Therefore, we compose them in an open framework capable of fully reconstructing complex real-world scenes from a single image.
@@ 3 Method
Figure 3: Detailed Overview over the processing steps of each instance. An image is processed through the scene analysis part of our framework as described in Figure 2. Then, we add an object recognition information for diffusion guided completion to restore partially occluded objects. Lastly, we perform reconstruction and align the result back to the input view space.
@@ 3.1 Scene analysis
Our framework decouples the object-agnostic from the object-specific processing, pursuing a balance between the representational power and the generalization capabilities of the integrated modules: That is, the perspective properties, semantic labels, and depth information are best retrieved by perceiving the scene as a whole
The geometry of a scene is characterized by its layout and the amodal shape of the contained objects. The layout of a scene refers to the surfaces that enclose the space (e.g., walls) and the 3D locations of the objects. To model the layout, we unproject a monocular depth estimation $D$, of the input image using predicted camera calibration parameters, $K_{img}$, as a point cloud in the 3D view space, $P^{view}\in\mathbb{R}^{3}$, and adopt it as our guide for positioning the scene components. A 2.5D representation is not sufficient to fully describe the layout of a scene as it only provides information for the visible parts. Still, it can be used to integrate individually reconstructed 3D objects into the scene (Section 3.2) and for background estimation (Section 3.3).
To parse the image, we opt for an entity segmentation [60, 59] approach in contrast to conventional 2D object detection, which enables us to segment all semantically-meaningful entities without being constrained to a predefined set of classes. This step partitions the image $I$ into instances $\{M_{i}\}$ that can be individually reconstructed to compose the whole scene. Furthermore, we consider the natural separation of the instances in thing and stuff using a universal image segmentation model. This facilitates our method to tailor the reconstruction process for each group, leveraging their unique properties (objects vs background). In this stage, we also label the identified entities, which can provide more context for instance processing.
@@ 3.2 Instance processing
Let an object $O_{i}$ be a well-defined shape categorised as thing, identified by its entity mask $M_{i}$, with corresponding RGB-D region $I_{i}=I[M_{i}]$, $D_{i}=D[M_{i}]$, and a label $L_{i}$. The processing steps are illustrated in Figure 3. We reconstruct each instance individually to fully benefit from a view-conditioned 2D diffusion model $\mathcal{Z}$ [48, 14], trained using multi-view images of mostly single objects from large scale collections [13, 14]. As the images used to train these models are rendered with a fixed predefined camera configuration $K_{crop}$, they generalize poorly to in-the-wild object crops. Therefore, we propose to address the domain-shift via reprojection of the object-associated pixels. To this end, we identify the virtual camera that together with the desired intrinsics $K_{crop}$ closely matches the observed image. Then, we use the transformation to project the unprojected pixels $P_{i}^{view}$ to a crop $C_{i}$ that resembles the training domain of $\mathcal{Z}$.
@@
For simple scenes in which there is no occlusion between instances, the crop $C_{i}$ represents the full view of the object $O_{i}$. However, this is not the case for most real-world scenes, and directly feeding $C_{i}$ to the object reconstruction method $\mathcal{R}$ would result in an incomplete object. Therefore, we propose to recover the missing parts of $C_{i}$ by leveraging the image prior embodied by pre-trained large-scale diffusion models like Stable Diffusion. We approach the task named amodal completion which deals with recovering the shape and appearance of partially visible instances as an image-to-image translation problem. Specifically, we train a model to predict the view $\hat{C}_{i}$ of the full object $O_{i}$ conditioned on the object parts depicted in $C_{i}$ and the label $L_{i}$. Given the difficulty of collecting well-segmented training images guaranteed to contain complete objects, we generate a synthetic dataset specifically for this task. We render synthetic objects from a large collection and then obtain the conditioning images by masking out parts of the object with a randomly overlaying silhouette of another object. For more details about the dataset generation and differences between amodal completion and inpainting please see the supplementary material. We use this dataset to fine-tune Stable Diffusion, following the methodology of InstructPix2Pix [6], and concatenate the encoded conditioning image to the noisy latent. The weights added to the base network are initialized with zero, while the rest are taken from the pre-trained model to benefit from the learned image prior.
@@
The complete crop $\hat{C}_{i}$ can now be used as input for the single-image 3D object reconstruction method $\mathcal{R}$. Using a view-conditioned diffusion prior enables the model to reconstruct a wide range of objects from a single view without the need for 3D training supervision. The object is reconstructed using a differentiable 3D representation (e.g., neural fields [52, 53] or 3D Gaussians [35]) that is either directly fitted to multi-view images generated by the diffusion prior, or by optimizing a Score Distillation Sampling-based loss [58] against the diffusion prior. Then, a polygonal mesh $R_{i}^{obj}$ aligned with the input crop $\hat{C}_{i}$ is extracted from the 3D representation using Marching Cubes [49, 70]. The obtained mesh can optionally be further fine-tuned to refine the texture of the reconstructed object.
@@
We transform the reconstructed instance from the object space (DreamGaussian) $R_{i}^{obj}$ to the view space (our scene) with the inverse transformation determined by the virtual camera used for projection. The obtained mesh $R_{i}^{view}$ is aligned to the object points $P_{i}$ up to an unknown scale factor $s_{i}$; this is because $\mathcal{R}$ reconstructs objects at an arbitrary scale. We estimate $s_{i}$ as the scale factor that minimizes the distance between the visible points in $R_{i}^{view}$ and $P_{i}^{view}$.